40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 7 18:20:01 2011, kiwamu, Summary, Green Locking, plans 
    Reply  Tue Mar 8 15:33:09 2011, kiwamu, Summary, Green Locking, plan on Mar.8th 
Message ID: 4382     Entry time: Mon Mar 7 18:20:01 2011     Reply to this: 4387
Author: kiwamu 
Type: Summary 
Category: Green Locking 
Subject: plans 
This week's goal is to investigate the source of the differential noise and to lower it.
 
Plans for tonight
 - realign GREEN_TRANS PD at the PSL table
 - update the noise budget
 - take spectrum of the differential noise
 - investigate a noise coupling to the differential noise especially from the intensity noise
 - update the noise budget again
 
Plans for this week :
 - Auto alignment scripts for green (Kiwamu)
 - connect the end REFL_DC  to an ADC (Kiwamu)
 - make an active phase rotation circuit for the end PDH (undergrads)
 - bounce-roll notches (Suresh)
 - optimization of the suspensions including the input matrices and the Q-values (Jenne)
 - optimization of MFSS (Koji/Rana/Larisa)
 - rewire the mechanical shutter on the 1X9 binary outputs (Steve)

 

ELOG V3.1.3-